Q4) ) You are given the following marginal utilities of goods X and Y obtained by a consumer. Given that price of X = Rs. 2.5, price of Y = Rs. 1 and income = Rs. 11, find out the optimal combination of goods

No. of unit consumed of a commodity: 1, 2, 3, 4, 5, 6, 7

Marginal utilities of X: 10, 12.5, 10, 7.5, 5, 2.5, 0.5

Marginal utilities of Y: 10, 9, 8, 7, 6, 5, 4



1
Expert's answer
2020-07-30T11:10:36-0400
"p_x=2.5"

"p_y=1"

"I=11"

"2.5x+y=11"

"\frac {MU_x}{p_x}= \frac {MU_y}{p_y}"


"MU_x=2.5MU_y"

"x=2, y=6"